OpenStudy (anonymous):

If a rectangle has a perimeter of 70, a length of x and a width of x - 9, find the value of the length of the rectangle..can someone help please?

OpenStudy (anonymous):

70=2(x+x-9)

OpenStudy (anonymous):

2x-9=35 x=22

OpenStudy (anonymous):

length 22, width 13

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

perimeter= 2(L+B)

OpenStudy (anonymous):

L+B?

OpenStudy (anonymous):

LENGTH + WIDTH

OpenStudy (anonymous):

okay...Thanks!